### 14 Dec 2023 v1

@@ 1 Introduction
Graph Neural Networks (GNNs) have gained popularity as a neural network-based approach for handling graph-structured data, leveraging their capacity to merge two information sources through the propagation and aggregation of node feature encodings along the network’s connectivity [14]. Nodes within a network can be organized into communities based on similarities in associated features and/or edge density [29]. Analyzing the network structure to identify clusters or communities of nodes proves valuable in addressing real-world issues like misinformation detection [22], genomic feature discovery [3], and social network or research recommendation [41]. As an unsupervised task, the identification of node clusters relies on latent patterns within the dataset rather than on “ground-truth” labels. Clustering holds significance for emerging applications lacking associated ground truth. Evaluating performance in discovering unknown information becomes crucial for applications where label access is restricted. Many graph applications involve millions of nodes, and datasets mimicking realistic scenarios exhibit low labeling rates [11].
@@ 2.1 Problem Definition
The problem definition of community detection on attributed graphs is defined as follows. The graph, where $N$ is the number of nodes in the graph, is represented as $G=(A,X)$, with the relational information of nodes modelled by the adjacency matrix $A\in\mathbb{R}^{N\times N}$. Given a set of nodes $V$ and a set of edges $E$, let $e_{i,j}=(v_{i},v_{j})\in E$ denote the edge that points from $v_{j}$ to $v_{i}$. The graph is considered weighted so, the adjacency matrix $0<A_{i,j}\leq 1$ if $e_{i,j}\in E$ and $A_{i,j}=0$ if $e_{i,j}\notin E$. Also given is a set of node features $X\in\mathbb{R}^{N\times d}$, where $d$ represents the number of different node attributes (or feature dimensions). The objective is to partition the graph $G$ into $k$ clusters such that nodes in each partition, or cluster, generally have similar structure and feature values. The only information typically given to the algorithms at training time is the number of clusters $k$ to partition the graph into. Hard clustering is assumed, where each community detection algorithm must assign each node a single community to which it belongs, such that $P\in\mathbb{R}^{N}$ and we evaluate the clusters associated with each node using the labels given with each dataset such that $L\in\mathbb{R}^{N}$.
@@ 2.2 Models
We consider a representative suite of GNNs, selected based on factors such as code availability and re-implementation time. In addition to explicit community detection algorithms, we also consider those that can learn an unsupervised representation of data as there is previous research that applies vector-based clustering algorithms to the representations [21].
@@ 2.3 Hyperparameter Optimisation Procedure
2.4 Suite of Tests
Table 2: The datasets and associated statistics.

### 04 Jan 2024 v2

@@ 1 Introduction
Graph Neural Networks (GNNs) have gained popularity as a neural network-based approach for handling graph-structured data, leveraging their capacity to merge two information sources through the propagation and aggregation of node feature encodings along the network’s connectivity [17]. Nodes within a network can be organized into communities based on similarities in associated features and/or edge density [33]. Analyzing the network structure to identify clusters or communities of nodes proves valuable in addressing real-world issues like misinformation detection [25], genomic feature discovery [4], and social network or research recommendation [49]. We consider unsupervised neural approaches to community detection that do not use any ground truth or labels during training to optimise the loss functions. As an unsupervised task, the identification of node clusters relies on latent patterns within the dataset rather than on “ground-truth” labels. Clustering holds significance for emerging applications lacking associated ground truth. Evaluating performance in discovering unknown information becomes crucial for applications where label access is restricted. Many graph applications involve millions of nodes, and datasets mimicking realistic scenarios exhibit low labeling rates [14].
@@ 1.1 Related Work
Su et al. [36] provides a taxonomy of existing node clustering methods, comparing representative works in the GNN space, as well as deep non-negative matrix factorisation and sparse filtering-based methods but do not empirically evaluate the methods discussed. Chunaev [5] provides an overview of community detection methods but omit mention of GNN methods. Tu et al. [41] provides an overview of network representation learning for community detection and proposes a novel method to incorporate node connectivity and features, but does not compare GNN methods in the study. Ezugwu et al. [9] surveys clustering algorithms providing an in-depth discussion of all applications of clustering algorithms, but does not includes an investigation into methods that perform clustering on graphs. There are many traditional community detection methods that are not based on neural networks. Louvain [3] is a classic community detection algorithm that uses two phases: in the modularity optimisation phase, nodes are randomly ordered then added and removed from communities until there is no significant change in modularity; nodes with the same community are collected and represented as a single node in community aggregation. Leiden [39] is an extension of Louvain, designed to fix the tendency to discover weakly connected communities. Leiden has some subtle differences which include the addition of a refinement of partitions phase in-between the modularity optimisation and community aggregation phases of Louvain. Rather than the greedy merging of Louvain based on the largest increase of modularity, the chance of merging increases in proportion to the modularity. Also, Leiden uses a ‘fast local move procedure’, whereby it only visits the nodes who’s neighbourhoods have been changed in the last iteration of the modularity optimisation. Label Propagation [31] uses network structure to find communities of densely connected nodes. This works by initialising every node with a unique label and at every iteration, each node is reassigned the label that the majority of it’s neighbours have. These traditional methods are transductive [43] as they do not learn a function that can be applied to unseen data, whereas in this work we only consider inductive GNN methods which produce a model that can can predict community assignments for data that it has not been trained on.
@@ 2.1 Problem Definition
The problem definition of community detection on attributed graphs is defined as follows. The graph, where $N$ is the number of nodes in the graph, is represented as $G=(A,X)$, with the relational information of nodes modelled by the adjacency matrix $A\in\mathbb{R}^{N\times N}$. Given a set of nodes $V$ and a set of edges $E$, let $e_{i,j}=(v_{i},v_{j})\in E$ denote the edge that points from $v_{j}$ to $v_{i}$. The graph is considered weighted so, the adjacency matrix $0<A_{i,j}\leq 1$ if $e_{i,j}\in E$ and $A_{i,j}=0$ if $e_{i,j}\notin E$. Also given is a set of node features $X\in\mathbb{R}^{N\times d}$, where $d$ represents the number of different node attributes (or feature dimensions). The objective is to partition the graph $G$ into $k$ clusters such that nodes in each partition, or cluster, generally have similar structure and feature values. The only information typically given to the algorithms at training time is the number of clusters $k$ to partition the graph into. Hard clustering is assumed, where each community detection algorithm must assign each node a single community to which it belongs, such that $P\in\mathbb{R}^{N}$ and we evaluate the clusters associated with each node using the labels given with each dataset such that $L\in\mathbb{R}^{N}$. Metrics that compare to the ground truth labels are allowed to be used for early-stopping and for hyperparameter optimisation.
@@ 2.2 Models
We consider a representative suite of GNNs, selected based on factors such as code availability and re-implementation time. In addition to explicit community detection algorithms, we also consider those that can learn an unsupervised representation of data as there is previous research that applies vector-based clustering algorithms to the representations [24]. The following are GNN architectures that learn representations of attributed graphs without a comparison to any labels during the training process. All these methods have hyperparameters which control trade-offs in optimisation.
@@ 2.3 Hyperparameter Optimisation Procedure
Table 2: The datasets and associated statistics.
2.4 Suite of Tests